OpenStudy (anonymous):

26 + (–3) ∙ (–8) please help me with 4 others as well

OpenStudy (anonymous):

i think its -184

OpenStudy (anonymous):

A. –184 B. –2 C. 2 D. 50

OpenStudy (alexandervonhumboldt2):

-*-=+ first do multiplication, then add (-3)*(-8)=3*8=24 24+26=50

OpenStudy (anonymous):

hu?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

can you help me?

alones (alones):

Just simplify it's 50

OpenStudy (anonymous):

ah ok got time for more?

alones (alones):

of coarse that';s up to you...

OpenStudy (anonymous):

What is the value of 3 • (–8) – 2 • (–5)?

alones (alones):

Simplify the expression and you'll get -14 do i need to explain???:) maybe it takes you're time to read

OpenStudy (anonymous):

no i have mentle memory problems

alones (alones):

really or your just joking well that's fine all of us are not perfect o/k what's your next Q?

OpenStudy (anonymous):

Simplify. –12 – 4.8 ∙ (–2 + –8) A. –168 B. 36 C. 72 D. 168

alones (alones):

o.k so it's

alones (alones):

36

OpenStudy (anonymous):

thx heres the next

alones (alones):

sure i'm there can you answer ma question too at the end??

OpenStudy (anonymous):

Which expressions, when evaluated with y = –7, are positive in value? Choose all answers that are correct. A. 18 + 2y B. –3y + (–17) C. –2y – 23 D. –30 – 4y ill try

alones (alones):

This one is pretty tuff but still i'll explain so If you replace y with -7 you get: a. 18 +2x-7 b. -3x-7 + (-17) c. (-2x-7)-23 d. -30-4x-7 you get: a. 18-14 b. 21-17 c. 14-23 d. -30+28 this solves to: a. 4 b. 4 c. -9 d. -2 so the positive answers are a and b

OpenStudy (anonymous):

thx heres the last one

alones (alones):

o.k i'm ready

OpenStudy (anonymous):

Evaluate the expression when t = –6.4 and u = –2.8. t – 4u A. –17.6 B. 4.8 C. 17.6 D. 29.12

alones (alones):

wait i'm confused is there more options or only one??

OpenStudy (anonymous):

single answer

alones (alones):

o.k that gives me alot

OpenStudy (anonymous):

realy or sarcasem?

OpenStudy (anonymous):

wait what grade ya?

alones (alones):

wait i need to solve first i'l answer them in a momment ..

OpenStudy (anonymous):

i looked it up and it says 4.8

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

you afk?

alones (alones):

I can't find the right answer i got -13.2 But the first time i got a yeah probably that one cause i just got 4.8

alones (alones):

sorry i don't understand what "afk" stnds for??

OpenStudy (anonymous):

ill tell ya the scor

OpenStudy (anonymous):

100%!!!!!!!!!!!!!!

alones (alones):

o.k i'll be glad to hear it so what is afk

OpenStudy (anonymous):

it means away from keyboard

alones (alones):

I' m so excited for your score so your asking me if i'm away from keyboard right now i'm see ya maybe we can speak after

OpenStudy (anonymous):

sorry i dint get to helping you because i love helping people

OpenStudy (anonymous):

lol im male

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

its fine looking for freinds?

OpenStudy (anonymous):

id be happy privet chatting ya my cell its np i had a bad time with low freinds so id be happy to

OpenStudy (anonymous):

im not dont worry sending cell num now

OpenStudy (anonymous):

got skype then?

OpenStudy (anonymous):

kk then see ya round

OpenStudy (anonymous):

iowa

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

closing q